OpenStudy (anonymous):

suppose you pay $24 for a pair of shoes that have been discounted 20%. what is the original price of the shoes

OpenStudy (anonymous):

It would be $30 20% is 1/5 30/5= 6 30-6=24

OpenStudy (goten77):

well lets say it was 50% off.. that means you only payed half so if its 20% off that means you only payed 80% of the original but to do a problem like this its best to set up a proportion 24/x = 4/5 the 4/5 is how much was actually payed since 1/5 was taken away... then solve for x 24= 4x/5 120=4x 30=x